OpenStudy (anonymous):

Solve the proportion. 6/a = 18/27

pooja195 (pooja195):

Use cross multiplication \[\huge~\rm~\frac{ 6 }{ a }=\frac{ 18 }{ 27}\]

OpenStudy (anonymous):

divide 18/27 by 3

OpenStudy (anonymous):

When you cross multiply you can find the answer a lot more quicker (in my opinion). (6 times 27) = (a times 18)

OpenStudy (anonymous):

From there you can divide 18 on both sides of the equation to get you answer.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Use cross multiplication \[\huge~\rm~\frac{ 6 }{ a }=\frac{ 18 }{ 27}\] \(\color{#0cbb34}{\text{End of Quote}}\)